# M3LEO: A Multi-Modal, Multi-Label Earth Observation Dataset Integrating Interferometric SAR and RGB Data

### v1

@@ Abstract
Satellite-based remote sensing has revolutionised the way we address global challenges in a rapidly evolving world. Huge quantities of Earth Observation (EO) data are generated by satellite sensors daily, but processing these large datasets for use in ML pipelines is technically and computationally challenging. Specifically, different types of EO data are often hosted on a variety of platforms, with differing degrees of availability for Python preprocessing tools. In addition, spatial alignment across data sources and data tiling for easier handling can present significant technical hurdles for novice users. While some preprocessed Earth observation datasets exist, their content is often limited to optical or near-optical wavelength data, which is ineffective at night or in adverse weather conditions. Synthetic Aperture Radar (SAR), an active sensing technique based on microwave length radiation, offers a viable alternative. However, the application of machine learning to SAR has been limited due to a lack of ML-ready data and pipelines, particularly for the full diversity of SAR data, including polarimetry, coherence and interferometry. In this work, we introduce M3LEO, a multi-modal, multi-label Earth observation dataset that includes polarimetric, interferometric, and coherence SAR data derived from Sentinel-1, alongside Sentinel-2 RGB imagery and a suite of labelled tasks for model evaluation. M3LEO spans 17.5TB and contains approximately 10M data chips, each measuring 4x4 km, across six diverse geographic regions. The dataset is complemented by a flexible PyTorch Lightning framework, with configuration management using Hydra, to accommodate its use across diverse ML applications in Earth observation. Additionally, we provide tools to process any dataset available on popular platforms such as Google Earth Engine for seamless integration with our framework. Initial experiments validate the utility of our data and framework, and show that SAR imagery contains information additional to that extractable from RGB data alone. Data is available at huggingface.co/M3LEO, and code at github.com/spaceml-org/M3LEO.
@@ 1 Introduction
Satellite-based Earth observation data is fundamental in addressing global problems in a rapidly changing world, offering large-scale, high resolution, high frequency data for applications from tracking wildfires [1] and deforestation [2] to refugee settlement mapping [3] and war zone damage assessment [4][5]. Information from these tasks is critical in crafting responses to man-made [6] and environmental crises [7], but is constrained by the use of optical (wavelengths from visible to near-infrared) sensing data. Such sensors are unable to operate in adverse weather or cloudy conditions [8], or at night, limiting their usefulness for time-critical tasks such as natural disaster management [7], environmental protection [9] or maritime surveillance [10]. Synthetic Aperture Radar (SAR) sensing presents an alternative that is able to overcome these limitations.
@@
Unlike optical sensors, SAR instruments actively illuminate terrain using microwave pulses, ensuring visibility without the need for daylight. These long-wavelength pulses can penetrate cloud cover and other adverse atmospheric conditions such as dust, making SAR-based sensing a valuable alternative to optical sensors for robust day-night coverage. Additionally, microwave radiation can penetrate some solid objects such as small-scale vegetation or soils - allowing, for example, measurements of properties relating to the soil moisture under vegetation [11] or the identification of archaeological features hidden below ground [12]. In addition to exploiting the wavelength and active illumination of SAR data, the complex nature of SAR signals - returning both amplitude and phase - can also be leveraged to provide insights beyond what is possible using optical data. Coherence, the complex correlation between pairs of SAR acquisitions, has been used successfully for tasks including flood detection [13][14], detection of urban damage [15] and forest canopy height measurement [16]. The phase difference between co-registered SAR acquisitions, measured through interferometry, enables the detection of surface height changes with millimetre accuracy, independently of the horizontal resolution of the sensor. This capability is critical for monitoring geological phenomena such as earthquakes [17], landslides [18], glacial movement [19][20] and infrastructure deformation [21].
@@
While SAR offers opportunities to overcome the limitations of optical sensors, and to give insights that are impossible to provide using visible wavelengths, it is associated with substantial additional complexity. The automated analysis of optical data, sometimes used in fusion with SAR data, has seen great success in recent years [22][23][24] - including the development of large foundation models able to make use of planetary-scale datasets [25][26]. The application of large-scale deep learning to SAR data without simultaneous use of optical data, however, is limited [27]. The complexities of processing SAR data, particularly estimating coherence and performing interferometry, mean that the full diversity of SAR data types is not available at scale in formats compatible with machine learning (ML) pipelines. Consequently, large-scale work applying deep learning to SAR data is limited, especially for the full diversity of available SAR datatypes.
@@
To address these challenges we introduce M3LEO, a large-scale, multi-modal, multi-label Earth Observation dataset comprising polarimetric, interferometric and coherence data for SAR as well as Sentinel-2 RGB data and labelled data for several tasks. We provide this data pre-processed as ML-readable, tiled images to abstract complex parameter choices typically made by domain experts. We also include a flexible PyTorch Lightning framework, parameterised using Hydra [28], to further minimise barriers to usage. We perform a small set of baseline experiments on a subset of our data to illustrate the usefulness of this framework, and the usefulness of SAR data for Earth observation problems. We demonstrate that this data can be used in tandem with more common RGB data with a set of experiments using models with inputs from multiple data sources. Finally, in addition to the pre-formatted data we offer for download, we provide tools enabling ML practitioners to process any data retrievable from Google Earth Engine into the same tiled format, such that it can be used in our framework.
@@ 2 Related Work
Deep learning has been applied over the last decade to curated optical imagery with great success [29][30][31][32], including the recent development of large, self-supervised foundation models [33][34][35]. Such models have been extraordinarily successful in tasks such as semantic segmentation [36][37], image classification [32], [36] [38] and object detection [39][36]. EO data from optical sensors has similarly been the subject of success for deep learning practitioners. Early work focusing on small, fully supervised models, showed great promise in a huge range of tasks, including land cover classification [22], biomass measurement [23], road detection [40][24] and flood mapping [41], although many models have been limited in scope to a small geographic area [42][23][43][44]. More recent work focuses on the development of large foundation models, often self supervised [27][45], which are readily adaptable to a range of downstream tasks and geographic areas [26].
@@
Work applying foundation models to optical EO data has been enabled by the wealth of easily accessible open data. As well as being available as raw products from satellite data providers such as ESA [46], many datasets comprising optical satellite imagery in ML-ready formats exist [47][48][49][50], across a range of spatial resolutions [51][52], and for multiple time-points [48][53], although may be limited in other scopes - for example, not having aligned task labels [54][53] or being limited to a single region [48]. We provide data at a large scale (14.1% of the surface of the Earth) with aligned labels for a diverse set of tasks.
@@
The application of deep learning to SAR data is less comprehensive. A body of work applying deep learning directly to SAR data exists [55][56][57][58] [59], but data limitations mean that geographic or temporal generalisability, often lacking in remote sensing models [60], has not clearly been shown. The development of foundation models for SAR may prove to be productive in obtaining geographic and temporal generalisability. To create foundation models, SAR is commonly applied alongside optical imagery in data fusion-based approaches, at small geographic scales [61][62], but with little attribution regarding whether such approaches can work well without optical data. Some work exists exploiting schemes such as masked autoencoding [63][64], contrastive learning [65], or knowledge distillation [66][67] to develop foundation models for polarimetric SAR - and shows that strong geographic generalisabilty is obtainable when using SAR data at large scales [63][67]. With M3LEO, we make polarimetric, coherence and interferometric SAR data available at such scales for the development of SAR foundation models.
@@
The full diversity of SAR datatypes, such as interferometry, coherence and backscatter, have seen little application in machine learning but have enjoyed a rich tapestry of applications in other contexts. Interferometry, for example, is often used to track earthquakes [17], landslides [18] and glacial movement [19][20], in a manner that is both more repeatable - being immune to adverse weather conditions - and more accurate - being able to track millimetre-scale height changes - than methods using optical data. Coherence has been used with success in urban damage assessment [15], flood detection [13][14] and canopy height measurement [16]. We make these datatypes available at a multi-continental scale, alongside both polarimetric SAR and optical data, in M3LEO.
3 Dataset
@@ 3.1 SAR Datasets
The many benefits of SAR data are met increased complexity compared to optical data. SAR sensors are active - that is, they emit microwave pulses (5.6cm wavelength for Sentinel-1) and measure backscatter rather than imaging the Earth under passive illumination from the Sun. This enables day-night operation and the penetration of atmospheric obstructions such as cloud and dust. Unlike sunlight, the pulses emitted by SAR sensors are polarised, with the ability to emit pulses polarised either horizontally or vertically with respect to the Earth. SAR sensors are also able to measure the polarization of the backscatter, and capture both amplitude and phase. This complex signal allows studying the geometry of surface-level objects, in addition to their reflectances. As an example, higher amplitudes are measured when surface features align with the polarisation of the emitted pulse. As with optical imagery, objects smaller than the measurement wavelength are invisible to the sensor.
The use of SAR data is further complicated by the use side-looking radar. SAR sensors operate with the emitter and receiver aimed laterally, rather than vertically, as for optical sensors. Since radar operates by measuring the time of arrival for a backscattered signal, aiming the sensor vertically would make it impossible to distinguish between targets at an equal distance to the left or right of the direction of travel. This is corrected by side-looking, with the sensor aimed laterally such that the entire field of view is to one side of the satellite. Although side-looking corrects directional ambiguity, it necessitates complex post-processing to correct the resulting geometrical distortions and radiance redistribution [68][69][70], and results in the same terrain being imaged differently depending on the direction of satellite travel, as the terrain is illuminated from the opposite side. We provide data in both ascending (northwards) and descending (southwards) satellite directions in M3LEO.
We provide three products derived from SAR data in this dataset - polarimetric amplitude, intereferometry, and coherence. We give a brief background on each of these datatypes below, although we omit most technical details regarding their construction as they are beyond the scope of this work. References are provided for users who are interested in further background on these datatypes.
@@
Polarimetric amplitude measures the power of the backscattered signal received by the sensor. We provide amplitude data derived from ESA Sentinel-1 Level 1 Ground Range Detected SAR data (S1GRD), as available in Google Earth Engine (GEE) [^1] https://developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S1_GRD. Phase information is not provided via GEE and it is therefore impossible to produce further SAR datatypes from data available on the GEE platform. We provide data measuring vertically polarised and horizontally polarised returns from a vertically polarised emission (referred to as VV and VH respectively). Data is provided for imagery from both ascending and descending trajectories. We refer users to [71] for a more detailed breakdown of the theory behind polarimetric SAR data. S1GRD data is of 10m/pixel resolution and from 2020 only in M3LEO at the time of writing.
@@
Interformetry measures the phase difference between pairs of acquisitions over the same terrain. These phase differences provide data about small-scale displacements, which can be measured modulo the wavelength. Post-processed interferograms are unwrapped by computing accumulated modulo-wavelength displacements. As the scale at which these displacements can be measured depends on wavelength, rather than horizontal resolution, interferometric phase difference can be used to measure surface height changes with millimetre accuracy. We provide interferometric data computed using select pairs of Sentinel-1 acquisitions, from ASF ARIA Geocoded UNWrapped Inteferogram data (GUNW), as available in ASF vertex [^2] https://asf.alaska.edu/datasets/daac/aria-sentinel-1-geocoded-unwrapped-interferograms/[72]. We include all available acquisition pairs with time deltas of less than 72 days. We refer users to [73] for a detailed treatment of the processing steps required to construct interferograms in addition to the underlying physics. GUNW data from 2020 is provided exclusively at the time of writing, at a resolution of approximately 90m/pixel.
@@
To provide meaningful coherence values, expectations are computed within a small spatial window surrounding each pixel. The resulting resolution of coherence maps is therefore lower than that of the original acquisitions. Man-made structures typically exhibit high coherence, as they are stable across acquisitions. Forests and other vegetation larger than the wavelength of the instrument have lower coherence. Coherence is affected in all cases by additional decorrelation factors not relating directly to terrain, such as doppler centroid difference or thermal noise. [74] and [75] provide more detailed treatments of coherence estimation. We provide coherence estimates from the Global Seasonal Sentinel-1 Interferometric Coherence (GSSIC) dataset [76], using date-pairs with time deltas of 12, 24, 36 and 48 days, with one date-pair per tile per season. GSSIC data is of approximately 90m/pixel resolution and from 2020 only.
For both interferometry and coherence, the selection of acquisition pairs is critical. The acquisition pair selection process introduces a significant combinatorial challenge to managing SAR datasets - if every possible combination between all acquisitions were considered, the number of interferograms or coherence estimates would grow quadratically with the number of acquisitions. We pre-empt this issue by the provision of pre-selected date-pairs.
RGB
We additionally provide Red, Green and Blue (S2RGB) data sourced from the ESA Sentinel-2 mission [^3] https://developers.google.com/earth-engine/datasets/catalog/sentinel-2[46], as available in Google Earth Engine. We summarize this data as means per-month, and provide data for 2020 only at the time of writing.
3.2 Labelled Datasets & Digital Elevation Models

### v2

# M3LEO: A Multi-Modal, Multi-Label Earth Observation Dataset Integrating Interferometric SAR and Multispectral Data

@@ Abstract
Satellite-based remote sensing has revolutionised the way we address global challenges in a rapidly evolving world. Huge quantities of Earth Observation (EO) data are generated by satellite sensors daily, but processing these large datasets for use in ML pipelines is technically and computationally challenging. Specifically, different types of EO data are often hosted on a variety of platforms, with differing degrees of availability for Python preprocessing tools. In addition, spatial alignment across data sources and data tiling for easier handling can present significant technical hurdles for novice users. While some preprocessed Earth observation datasets exist, their content is often limited to optical or near-optical wavelength data, which is ineffective at night or in adverse weather conditions. Synthetic Aperture Radar (SAR), an active sensing technique based on microwave length radiation, offers a viable alternative. However, the application of machine learning to SAR has been limited due to a lack of ML-ready data and pipelines, particularly for the full diversity of SAR data, including polarimetry, coherence and interferometry. In this work, we introduce M3LEO, a multi-modal, multi-label Earth observation dataset that includes polarimetric, interferometric, and coherence SAR data derived from Sentinel-1, alongside multispectral Sentinel-2 imagery and a suite of auxiliary data describing terrain properties such as land use. M3LEO spans approximately 17M data chips, each measuring 4$\times$4 km, across six diverse geographic regions. The dataset is complemented by a flexible PyTorch Lightning framework, with configuration management using Hydra, to accommodate its use across diverse ML applications in Earth observation. Additionally, we provide tools to process any dataset available on popular platforms such as Google Earth Engine for seamless integration with our framework. We show that the distribution shift in self-supervised embeddings is substantial across geographic regions, even when controlling for terrain properties. Data is available at huggingface.co/M3LEO, and code at github.com/spaceml-org/M3LEO.
@@ 1 Introduction
Satellite-based Earth observation data is fundamental in addressing global problems in a rapidly changing world, offering large-scale, high resolution, high frequency data for applications from tracking wildfires [1] and deforestation [2] to refugee settlement mapping [3] and war zone damage assessment [4, 5]. Information from these tasks is critical in crafting responses to man-made [6] and environmental crises [7], but is constrained by the use of optical (wavelengths from visible to near-infrared) sensing data. Such sensors are unable to operate in adverse weather or cloudy conditions [8], or at night, limiting their usefulness for time-critical tasks such as natural disaster management [7], environmental protection [9] or maritime surveillance [10]. Synthetic Aperture Radar (SAR) sensing presents an alternative that is able to overcome these limitations.
@@
Unlike optical sensors, SAR instruments actively illuminate terrain using microwave pulses, ensuring visibility without the need for daylight. These long-wavelength pulses can penetrate cloud cover and other adverse atmospheric conditions such as dust, making SAR-based sensing a valuable alternative to optical sensors for robust day-night coverage. Additionally, microwave radiation can penetrate some solid objects such as small-scale vegetation or soils — allowing, for example, measurements of properties relating to soil moisture under vegetation [11] or the identification of archaeological features hidden below ground [12]. In addition to exploiting the wavelength and active illumination of SAR data, the complex nature of SAR signals — returning both amplitude and phase — can also be leveraged to provide insights beyond what is possible using optical data. Coherence, the complex correlation between pairs of SAR acquisitions, has been used successfully for tasks including flood detection [13, 14], detection of urban damage [15] and forest canopy height measurement [16]. The phase difference between co-registered SAR acquisitions, measured through interferometry, enables the detection of surface height changes with millimetre accuracy, independently of the horizontal resolution of the sensor. This capability is critical for monitoring geological phenomena such as earthquakes [17], landslides [18], glacial movement [19, 20], magma chamber growth [21] and infrastructure deformation [22].
@@
While SAR offers opportunities to overcome the limitations of optical sensors, and to give insights that are impossible to provide using visible wavelengths, it is associated with substantial additional complexity. The automated analysis of optical data, sometimes used in fusion with SAR data, has seen great success in recent years [23, 24, 25] — including the development of large foundation models able to make use of planetary-scale datasets [26, 27]. The application of large-scale deep learning to SAR data without simultaneous use of optical data, however, is more limited [28]. The complexities of processing SAR data, particularly estimating coherence and performing interferometry — which require processing phase information as using complex numbers — mean that the full diversity of SAR data types is not available at scale in formats compatible with machine learning (ML) pipelines.
@@
To address these challenges we introduce M3LEO, a large-scale multi-modal Earth Observation dataset comprising polarimetric, interferometric and coherence data for SAR as well as Sentinel-2 data and auxiliary datasets describing surface properties such as land use, biomass measurements and elevation. We provide this data pre-processed as ML-readable, tiled images to abstract complex parameter choices typically made by domain experts. We also include a flexible PyTorch Lightning framework, parameterised using Hydra [29], to further minimise barriers to usage. We include a preliminary analysis on distribution shift for terrain properties and the appearance of SAR data across geographic regions. Finally, in addition to the pre-formatted data we offer for download, we provide tools enabling ML practitioners to process any data retrievable from Google Earth Engine into the same tiled format, such that it can be used in our framework.
@@ 2 Related Work
Deep learning has been applied over the last decade to curated optical imagery with great success [30, 31, 32, 33], including the recent development of large, self-supervised foundation models [34, 35, 36, 37]. Such models have been extraordinarily successful in tasks such as semantic segmentation [38, 39], image classification [33], [38] [40] and object detection [41, 38]. EO data from optical sensors has similarly been the subject of success for deep learning practitioners. Early work focusing on small, fully supervised models showed great promise in a huge range of tasks, including land cover classification [23], biomass measurement [24], road detection [42, 25] and flood mapping [43], although many models were limited in scope to a small geographic area [44, 24, 45, 46]. More recent work has focused on the development of large foundation models, often self supervised [28, 47, 37], which are readily adaptable to a range of downstream tasks and geographic areas [27].
@@
Work applying these models to optical EO imagery has been enabled by the wealth of easily accessible open data. As well as being available as raw products from satellite data providers such as ESA [48], many datasets comprising optical satellite imagery in ML-ready formats exist [49, 50, 51, 52], across a range of spatial resolutions [53, 54], and for multiple time-points [50, 55], although they may be limited in other scopes — for example, not having aligned task labels [56, 55] or being limited to a single region [50]. We provide such data at a large scale (14.1% of the land surface of the Earth) with a diverse set of auxiliary labels.
@@
Beyond optical data, the application of deep learning to SAR data is less comprehensive. A body of work applying deep learning directly to SAR data exists [57, 58, 59, 60] [61], but data limitations mean that geographic or temporal generalisability, often lacking in remote sensing models [62], has not clearly been shown. The development of foundation models for SAR may prove to be productive in obtaining geographic and temporal generalisability. To create foundation models, SAR is commonly applied alongside optical imagery in data fusion-based approaches [63, 64], but with little attribution regarding whether such approaches can work well without optical data. Some work exists exploiting schemes such as masked autoencoding [65, 66], contrastive learning [67], or knowledge distillation [68, 69] to develop foundation models for polarimetric SAR — and shows that strong geographic generalisabilty is obtainable when using SAR data at large scales [65, 69]. Many datasets providing ML-readable polarimetric SAR (polSAR) data exist [51, 55, 70, 71, 72, 73, 74, 75], although most do not provide interferometric SAR (inSAR) data [74, 75].
@@
The full diversity of inSAR datatypes, such as interferometry and coherence, have seen a number of applications in machine learning, and a rich tapestry of applications in other contexts. Interferometry, for example, is often used to track earthquakes [17], landslides [18] and glacial movement [19, 20], in a manner that is both more repeatable — being immune to adverse weather conditions — and more accurate — being able to track millimetre-scale height changes — than methods using optical data. Coherence has been used with success in urban damage assessment [15], flood detection [13, 14] and canopy height measurement [16].
A small number of datasets exist making these datatypes available to deep learning users, many of which are focused on specific events, tasks or locations. Hephaestus [76], for example, contains 216K interferometric SAR (inSAR) patches localised to volcanoes annotated with various labels describing volcanic activity. ISSLIDE [77] contains inSAR data from the French alps describing slow landslides, and Pol-InSAR-Island [78] comprises inSAR data describing land cover on Baltrum, a Frisian island. UrbanSARFloods [79] contains Sentinel-1 interferometric coherence data from a diverse set of global locations, but is limited to specific events. S1SLC_CVDL [80] opts to provide complex-valued single-look SAR data rather than processed interferometric data, from three manually selected Sentinel-1 scenes containing major population centres. We make inSAR and coherence data available at a multi-continental scale, alongside both polarimetric SAR and optical data, in M3LEO.
2.1 Comparison to Existing Datasets
We provide a comparison between a number of popular large-scale Earth observation datasets, including M3LEO, in Table 2.1. We define tile to mean a fixed location or area on the surface of the Earth, and chip as the content of some data product over that tile. Unlike in many of these datasets, we provide acquisitions from different seasons within the same year for the same tile as channels, rather than separate chips. The described number of chips in Table 2.1 therefore appears relatively lower for M3LEO compared to, for example, SSL4EO-S12, for a fixed number of satellite acquisitions. We instead measure the number of timepoints in years, rounded up for part-years. Of the four datasets offering Sentinel-1 SAR data, only M3LEO offers data from multiple years, and only M3LEO offers inSAR data (although see Section 1 for a brief description of available task-specific or localised interferometric datasets).
Regarding spatial coverage, of the datasets listed in Table 2.1 M3LEO is most similar in scale to SatlasPretrain and SSL4EO-L, with these three datasets being significantly larger than the remainder. Of these three datasets, only M3LEO provides SAR data of any modality.
The temporal coverage of M3LEO sits between SatlasPretrain and SSL4EO-L, although the aggregate number of years does not give a full picture — in SatlasPretrain, some acquisitions are available for a wider range of years for specific events, and in SSL4EO-L different data products are sometimes collected for non-overlapping years, meaning much of the dataset is not a parallel corpus. In M3LEO, the primary satellite data from Sentinel-1 and Sentinel-2 is available for 2018-2020 for all tiles.
Several of these datasets contain auxiliary information in addition to satellite acquisitions. Land cover labels are common, included in SSL4EO-L, SEN12MS, BigEarthNet and SatlasPretrain [51, 56, 72, 73]. Additional auxiliary datasets are sometimes available — SSL4EO-L, for example, includes more detailed crop classification data. SatlasPretrain introduced a number of novel additional labelled datasets including building and road polygons. M3LEO currently includes 4 auxiliary datasets - Land cover, vegetation cover, aboveground biomass and Digital Elevation Models (DEMs).
Many datasets are pre-sampled within their selected AOIs prior to distribution. Some datasets sample based on a manually specified distribution (for example, SSL4EO-S12 and SSL4EO-L sample locations based on Gaussian distributions centred on large cities), and some randomly. We include all available tiles within our AOIs — partly with a view to increasing data volume, but also to enable further research on sampling schemes. Selecting a good sampling strategy to diversify actively illuminated radar data with phase information is not straightforward, and it doubtful whether sampling schemes developed for optical EO imagery would transfer well to this data. The auxiliary data included in M3LEO, such as elevation and land use, may be useful for constructing such sampling schemes.
[h] Existing Datasets. Summary of popular large-scale Earth observation pre-training datasets. Dataset SAR Years Num. Num. Tile Size Coverage Sampling Tiles Chips* km km2 SSL4EO-S12 [55] Y 1 251K 3M 2.64$\times$2.64 1.75×$\times 10^{6}$ Targeted SSL4EO-L [51] N 6 250K 5M 7.92$\times$7.92 1.57×$\times 10^{7}$ Targeted SEN12MS [70] Y 1 181K 542K 2.56$\times$2.56 1.18×$\times 10^{6}$ Random SeCo [71] N 2 200K 1M 2.65$\times$2.65 1.40×$\times 10^{6}$ Random BigEarthNet [72] N 1 590K 1.2M 1.20$\times$1.20 8.50×$\times 10^{5}$ Targeted SatlasPretrain [73] N 1** 856K N/A 5.12$\times$5.12 2.13×$\times 10^{7}$ None M3LEO Y 3 1.05M 17.2M 4.48$\times$4.48 2.11×$\times 10^{7}$ None
    Heuristic only — some work, such as SSL4EO-S12, considers acquisitions at the same location from different seasons to be a seperate data chip.
    Contains additional historical images from 2016-2021 that are relevant to dynamic events such as floods.
3 Dataset & Framework
@@ 3.1 SAR Datasets
The many benefits of SAR data are met with increased complexity compared to optical data. SAR sensors are active — that is, they emit microwave pulses (5.6 cm wavelength for Sentinel-1) and measure backscatter rather than imaging the Earth under passive illumination from the Sun. This enables day-night operation and the penetration of atmospheric obstructions such as cloud and dust. Unlike sunlight, the pulses emitted by SAR sensors are polarised, with the ability to emit pulses polarised either horizontally or vertically with respect to the Earth. SAR sensors are also able to measure the polarization of the backscatter, and capture both amplitude and phase. This signal, typically stored as a complex number, allows studying the geometry of surface-level objects, in addition to their reflectances. As an example, higher amplitudes are measured when surface features align with the polarisation of the emitted pulse. As with optical imagery, objects smaller than the measurement wavelength are invisible to the sensor.
The use of SAR data is further complicated by the use of side-looking radar. SAR sensors operate with the emitter and receiver aimed laterally, rather than vertically, as for optical sensors. Since radar operates by measuring the time of arrival for a backscattered signal, aiming the sensor vertically would make it impossible to distinguish between targets at an equal distance to the left or right of the direction of travel. This is corrected by side-looking, with the sensor aimed laterally such that the entire field of view is to one side of the satellite. Although side-looking corrects directional ambiguity, it necessitates complex post-processing to correct the resulting geometrical distortions and radiance redistribution [81, 82, 83], and results in the same terrain being imaged differently depending on the direction of satellite travel, as the terrain is illuminated from the opposite side. We provide data in both ascending (northwards) and descending (southwards) satellite directions in M3LEO.
We provide three products derived from SAR data in this dataset — polarimetric amplitude, intereferograms, and coherence. We give a brief background on each of these datatypes below, although we omit most technical details regarding their construction as they are beyond the scope of this work. References are provided for users who are interested in further background on these datatypes.
@@
Polarimetric amplitude measures the power of the backscattered signal received by the sensor. We provide amplitude data derived from ESA Sentinel-1 Level 1 Ground Range Detected SAR data (S1GRD), as available in Google Earth Engine (GEE)111developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S1_GRD. Phase information is not provided via GEE and it is therefore impossible to produce further SAR datatypes from data available on the GEE platform. We provide data measuring vertically polarised and horizontally polarised returns from a vertically polarised emission (referred to as VV and VH respectively). Data is provided for imagery from both ascending and descending trajectories. We refer users to [84] for a more detailed breakdown of the theory behind polarimetric SAR data. S1GRD data is of 10 m/pixel resolution and provided for 2018-2020 as four seasonal averages per-year.
@@
Interformetry measures the phase difference between pairs of acquisitions over the same terrain. These phase differences provide data about small-scale displacements, which can be measured modulo the wavelength. Post-processed interferograms are unwrapped by computing accumulated modulo-wavelength displacements. As the scale at which these displacements can be measured depends on wavelength, rather than horizontal resolution, interferometric phase difference can be used to measure surface height changes with millimetre accuracy. We provide interferometric data computed using select pairs of Sentinel-1 acquisitions, from ASF ARIA Geocoded UNWrapped Inteferogram data (GUNW), as available in ASF vertex222asf.alaska.edu/datasets/daac/aria-sentinel-1-geocoded-unwrapped-interferograms/[85]. We include all available acquisition pairs with time deltas of less than 72 days. We refer users to [86] for a detailed treatment of the processing steps required to construct interferograms in addition to the underlying physics. GUNW data is provided for 2020 at a resolution of approximately 90 m/pixel.
@@
To provide meaningful coherence values, expectations are computed within a small spatial window surrounding each pixel. The resulting resolution of coherence maps is therefore lower than that of the original acquisitions. Man-made structures typically exhibit high coherence, as they are stable across acquisitions. Forests and other vegetation larger than the wavelength of the instrument have lower coherence. Coherence is affected in all cases by additional decorrelation factors not relating directly to terrain, such as doppler centroid difference or thermal noise. [87] and [88] provide more detailed treatments of coherence estimation. We provide coherence estimates from the Global Seasonal Sentinel-1 Interferometric Coherence (GSSIC) dataset [89], using date-pairs with time deltas of 12, 24, 36 and 48 days, with one date-pair per tile per season. GSSIC data is of approximately 90 m/pixel resolution and from 2020. A decay model is included with the GSSIC coherence data.
For both interferometry and coherence, the selection of acquisition pairs is critical. The acquisition pair selection process introduces a significant combinatorial challenge to managing SAR datasets — if every possible combination between all acquisitions were considered, the number of interferograms or coherence estimates would grow quadratically with the number of acquisitions. We pre-empt this issue by the provision of pre-selected date-pairs.
Optical data
We additionally provide data (S2SRM) sourced from the ESA Sentinel-2 mission333developers.google.com/earth-engine/datasets/catalog/sentinel-2[48], as available in Google Earth Engine. Data is provided from the L2A product (surface reflectance). We do not include top-of-atmosphere L1C reflectances. We summarize this data as monthly means of cloud-free pixels. We include four monthly averages for each of 2018-2020 — March, June, September and December. We include all Sentinel-2 bands with a resolution of $10$ m/pixel (red, green, blue, NIR) or $20$ m/pixel (vegetation red edge, SWIR 11/12).
3.2 Auxiliary Datasets